OpenStudy (monkey*):

Given f(x)=x^2+7 and g(x)=x-4/x. Find (g o f)(-1). Hint: (g o f)(-1)=g(f(-1))

OpenStudy (freckles):

so have you tried finding f(-1) and then finding g(put result of f(-1) here)

OpenStudy (freckles):

f(-1) tells us to use f(x)=x^2+7 we know to use f(x)=x^2+7 because of the f in f(-1)

OpenStudy (freckles):

f(x)=x^2+7 we can see here the input is the thing in f( ) example: f(a)=a^2+7 f(a+1)=(a+1)^2+7 f(input)=(input)^2+7 another example; f(2)=2^2+7 can you find f(-1)=?

OpenStudy (monkey*):

I'm a little lost, sorry.

OpenStudy (freckles):

ok do you see that f(x)=x^2+7 and you are also to find f(-1)?

OpenStudy (freckles):

f(x)=x^2+7 since we want to find f(-1) we need to replace those x's with -1

OpenStudy (monkey*):

Yes.

OpenStudy (freckles):

so can you find f(-1)?

OpenStudy (freckles):



OpenStudy (freckles):

let me know when you have found f(-1)

OpenStudy (monkey*):

Would it be f(-1)-1^2+7?

OpenStudy (freckles):

well there should be an equal sign somewhere there

OpenStudy (freckles):



OpenStudy (freckles):

can you simplify (-1)^2+7?

OpenStudy (monkey*):

8?

OpenStudy (freckles):

yes f(-1)=(-1)^2+7=(-1)(-1)+7=1+7=8 so f(-1)=8 so f(-1) is 8 now we had to find g(f(-1)) we just found that inside thing the f(-1) thing to be 8 replace f(-1) with 8 you are now really wanting to find g(8)

OpenStudy (freckles):

g(8) tells us to use the expression named g

OpenStudy (freckles):

g looks likes g(x)=x-4/x

OpenStudy (freckles):

just replace the old input x with the new input 8

OpenStudy (freckles):

simplify and then you are done

OpenStudy (monkey*):

So, g(8)=x-4/x?

OpenStudy (freckles):

you have to replace all the x's with 8 not just one of the x's

OpenStudy (freckles):

whatever you replace the first x with you have to replace the other x's on the other side also

OpenStudy (monkey*):

g(8)=8-4/8?

OpenStudy (freckles):

yes

OpenStudy (monkey*):

So simplify 8-4/8?

OpenStudy (monkey*):

@freckles

OpenStudy (freckles):

yes

OpenStudy (freckles):

I don't think teachers like unsimplified answers :p

OpenStudy (monkey*):

7.5 is my answer?

OpenStudy (monkey*):

And no, they don't lol.

OpenStudy (monkey*):

Thank you so much for the help!

OpenStudy (freckles):

for me to check your answer I have to know one this is it \[g(x)=\frac{x-4}{x} \text{ or } g(x)=x-\frac{4}{x}\]

OpenStudy (monkey*):

First one.

OpenStudy (freckles):

so it is actually g(x)=(x-4)/x not g(x)=x-4/x so g(8)=(8-4)/8=4/8=1/2 or .5

OpenStudy (freckles):

\[f(x)=x^2+7 \\ g(x)=\frac{x-4}{x} \\ g(f(-1)) = g((-1)^2+7)=g(1+7)=g(8)=\frac{8-4}{8}=\frac{4}{8}=\frac{1}{2}\]

OpenStudy (monkey*):

So what would be the answer?

OpenStudy (freckles):

well we just got that g(f(-1))=1/2 and we were looking for the value of g(f(-1)) so we are done